OpenStudy (anonymous):

6+/-2 square root of 3 ------------------ 4

OpenStudy (anonymous):

\[6+/-2\sqrt{3}\div4\]

OpenStudy (anonymous):

Wait what? 6 divded by negative 2\[\sqrt{3}\] divided by 4?

OpenStudy (nowhereman):

I guess it should be \[\frac{6\pm 2\sqrt{3}}{4}\], but what is the question?

OpenStudy (anonymous):

\[6/4 = 3/2\] so therefore it's\[[3 + or - \sqrt{3}]\div 2\]

OpenStudy (nowhereman):

ingenious

OpenStudy (anonymous):

Nowhereman: how'd you write the equation for that fraction? I've not managed to figure out how to do it?

OpenStudy (nowhereman):

It's latex. You can use the formula editor by clicking on Equation, or just put the latex code inside \ [ and \ ]: \frac{6 \pm 2\sqrt{3}}{4}

OpenStudy (anonymous):

oh! awsome, thanks. I didn't realize this was latex. I've been learning the syntax for the equation editor so I don't have to click the button and find the symbols I want, but I didn't realize it was standard latex. cool.